Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).
Answer by nerdybill(7384)

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).
.
Since our new line is parallel to:
y = 7x + 2
it must have the same slope (7).
Plug the slope (7) along with the given point (1,-3) into the "point-slope" form:
y-y1 = m(x-x1)

y-(-3) = 7(x-1)
y+3 = 7x-7
y = 7x - 10 (this is what they're looking for)